# Remote State Estimation over Unreliable Channels with Unreliable Feedback: Fundamental Limits

### v1

@@ V Conclusions
×$\qquad\quad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1}$
@@
[…]
×$\qquad\quad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1}$
where in the equalities we used the facts that $\Psi^{k+1,M}_{\epsilon^{n}}(\tilde{e})=\bar{\Psi}^{M-k-1}_{\epsilon^{n}}(\tilde{e})$ for any Gaussian variable $\tilde{e}$ and a suitable selection of $\bar{\boldsymbol{\alpha}}_{0:M-k-1}$, and that $\nu_{k+1}$ is independent of $\boldsymbol{\sigma}_{0:k}$, and the Fubini’s theorem; and in the inequality we used the hypothesis $\Psi^{M-k-1}_{\epsilon^{n}}(\tilde{e})\geq\Psi^{M-k-1}_{\epsilon^{s}}(\tilde{e})$ for any Gaussian variable $\tilde{e}$. Therefore,
@@
Step 3. Now, in the final step of our proof, we show that $\Phi(\epsilon^{\star},\delta^{\star})\leq\Phi(\epsilon^{s},\delta^{o})$, where $\epsilon^{\star}$ is a special form of $\epsilon^{s}$ and $\delta^{\star}$ is the same as $\delta^{o}$. Observe that, by Lemma 1 and Lemmas 3 and 5, when $\epsilon^{s}$ is used, $\delta^{o}$ must satisfy the recursive equation
$\hat{x}_{k}=A_{k-1}\hat{x}_{k-1}+\delta_{k-1}\gamma_{k-1}A_{k-1}(\check{x}_{k-1}-\hat{x}_{k-1})$ (49)
@@
$V_{k}(\mathcal{I}^{e}_{k})=\min_{\delta_{k}\in\{0,1\}}\bigg{\{}\alpha_{k}\delta_{k}+\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k}]$
$\qquad\qquad\qquad\qquad+\operatorname{tr}Q_{k+1}+\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k}]\bigg{\}}$ (50)
for $k\in\mathbb{N}_{[0,N-1]}$ with initial condition $V_{N}(\mathcal{I}^{e}_{N})=0$, where we used the additivity property of $V_{k}(\mathcal{I}^{e}_{k})$ and the fact that $\operatorname{\mathsf{E}}[\hat{e}_{k+1}^{T}\hat{e}_{k+1}|\mathcal{I}^{e}_{k}]=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k}]+\operatorname{tr}Q_{k+1}$ according to the tower properties of conditional expectations. The minimizing scheduling policy is then obtained as
$\delta_{k}=\mathds{1}_{\chi_{k}-\alpha_{k}\geq 0}$ (51)
where $\chi_{k}=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}+V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=0]-\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}+V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=1]$. Next, we prove by induction that $V^{e}_{k}(\mathcal{I}^{e}_{k})$ is a symmetric function of $\breve{e}_{k}$ and a function $R_{k}$. Clearly, the claim holds for time $N$. We assume that the claim holds for time $k+1$ and we will show that it also holds for time $k$. Let us write $\chi_{k}=\chi^{1}_{k}+\chi^{2}_{k}$, where $\chi^{1}_{k}=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},\delta_{k}=0]-\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},\delta_{k}=1]$ and $\chi^{2}_{k}=\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=0]-\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=1]$. From the properties of expectations, we have the following relation
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$ $=\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}^{T}\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$
$\quad+\operatorname{tr}\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}.$
@@
$\tilde{e}_{k+1}$ $=(1-\delta_{k}\gamma_{k})A_{k}\tilde{e}_{k}+K_{k+1}\nu_{k+1}.$
Accordingly, the conditional mean and the conditional covariance of $\tilde{e}_{k+1}$ given $\mathcal{I}^{e}_{k}$ and $\delta_{k}$ are obtained as
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$ $=(1-\lambda^{c}\delta_{k})A_{k}\breve{e}_{k}$
$\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$ $=(\lambda^{c}\delta_{k}-{\lambda^{c}}^{2}\delta_{k})A_{k}\breve{e}_{k}\breve{e}_{k}^{T}A_{k}^{T}$
$\quad+(1-\lambda^{c}\delta_{k})A_{k}R_{k}A_{k}^{T}$
$\quad+K_{k+1}N_{k+1}K_{k+1}^{T}$
@@
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}\big{|}\mathcal{I}^{e}_{k},\delta_{k}\Big{]}$ $=(1-\lambda^{c}\delta_{k})\breve{e}_{k}^{T}A_{k}^{T}A_{k}\breve{e}_{k}$
$\quad+(1-\lambda^{c}\delta_{k})\operatorname{tr}(A_{k}R_{k}A_{k}^{T})$
@@
This implies that $\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},\delta_{k}]$ is a symmetric function of $\breve{e}_{k}$ and a function of $R_{k}$, and that
@@
$\breve{e}_{k+1}$ $=(1-\bar{\theta}_{k})A_{k}\breve{e}_{k}+K_{k+1}\nu_{k+1}$ (52)
$R_{k+1}$ $=(1-\bar{\theta}_{k})A_{k}R_{k}A_{k}^{T}+\bar{\bar{\theta}}_{k}A_{k}\breve{e}_{k}\breve{e}_{k}^{T}A_{k}^{T}$ (53)
where $\bar{\theta}_{k}=\theta_{k}\delta_{k}\gamma_{k}+\lambda^{c}(1-\theta_{k})\delta_{k}$ and $\bar{\bar{\theta}}_{k}=(1-\theta_{k})(\lambda^{c}\delta_{k}-{\lambda^{c}}^{2}\delta_{k})$. By the hypothesis, $V^{e}_{k+1}(\breve{e}_{k+1},R_{k+1})$ is a symmetric function of $\breve{e}_{k+1}$ and a function of $R_{k+1}$. Plugging (52) and (53) into $V_{k+1}(\breve{e}_{k+1},R_{k+1})$, we can calculate $\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}]$ while all $\nu_{k+1}$, $\theta_{k}$, and $\gamma_{k}$ are averaged out. A straightforward consequence is that we can express $\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}]$ as a symmetric function of $\breve{e}_{k}$ and a function of $R_{k}$, since $\nu_{k+1}$ is a Gaussian variable with zero mean. Therefore, we deduce that $\chi^{2}_{k}$ is also a symmetric function of $\breve{e}_{k}$ and a function $R_{k}$. Thus, using (51), we conclude that the claim holds. ∎

### v2

@@ VIII Conclusions
$\qquad\quad\times\operatorname{\mathsf{p}}_{\epsilon^{n}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1}$
$=\int_{\mathbb{R}^{m}}\bar{\Psi}^{M-k-1}_{\epsilon^{n}}(K_{k+1}\nu_{k+1})\operatorname{\mathsf{p}}(\nu_{k+1})d\nu_{k+1}$
$\geq\int_{\mathbb{R}^{m}}\bar{\Psi}^{M-k-1}_{\epsilon^{s}}(K_{k+1}\nu_{k+1})\operatorname{\mathsf{p}}(\nu_{k+1})d\nu_{k+1}$
$\qquad\quad\times\operatorname{\mathsf{p}}_{\epsilon^{s}}(\boldsymbol{\nu}_{0:k+1}|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1)d\boldsymbol{\nu}_{0:k+1},$
@@
Step 3. Now, in the final step of our proof, we show that $\Phi(\epsilon^{\star},\delta^{\star})\leq\Phi(\epsilon^{s},\delta^{o})$, where $\epsilon^{\star}$ is a special form of $\epsilon^{s}$ and $\delta^{\star}$ is the same as $\delta^{o}$. Observe that, by Lemma 1 and Lemmas 3 and 5, for any $\epsilon^{s}$ that is adopted, $\delta^{o}$ must satisfy the recursive equation
$\hat{x}_{k}=A_{k-1}\hat{x}_{k-1}+u_{k-1}\gamma_{k-1}A_{k-1}(\check{x}_{k-1}-\hat{x}_{k-1}),$ (86)
@@
$V_{k}(\mathcal{I}^{e}_{k})=\min_{u_{k}\in\{0,1\}}\bigg\{\alpha_{k}u_{k}+\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k}]$
$\qquad\qquad\qquad\qquad+\operatorname{tr}Q_{k+1}+\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k}]\bigg\},$ (87)
for $k\in\mathbb{N}_{[0,N-1]}$ with initial condition $V_{N}(\mathcal{I}^{e}_{N})=0$, where we used the additivity property of $V_{k}(\mathcal{I}^{e}_{k})$ and the fact that $\operatorname{\mathsf{E}}[\hat{e}_{k+1}^{T}\hat{e}_{k+1}|\mathcal{I}^{e}_{k}]=\operatorname{\mathsf{E}}[\operatorname{\mathsf{E}}[\hat{e}_{k+1}^{T}\hat{e}_{k+1}|\boldsymbol{\gamma}_{0:k},\mathcal{I}^{e}_{k+1}]|\mathcal{I}^{e}_{k}]=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k}]+\operatorname{tr}Q_{k+1}$ according to the tower properties of conditional expectations. The minimising scheduling policy is then obtained as
$u_{k}=\mathds{1}_{\chi_{k}-\alpha_{k}\geq 0},$ (88)
where $\chi_{k}=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}+V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},u_{k}=0]-\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}+V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},u_{k}=1]$. Next, we prove by induction that $V_{k}(\mathcal{I}^{e}_{k})$ is a symmetric function of $\breve{e}_{k}$ and a function $R_{k}$. Clearly, the claim holds for time $N$. We assume that the claim holds for time $k+1$ and we will show that it also holds for time $k$. Let us write $\chi_{k}=\chi^{1}_{k}+\chi^{2}_{k}$, where $\chi^{1}_{k}=\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},u_{k}=0]-\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},u_{k}=1]$ and $\chi^{2}_{k}=\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},u_{k}=0]-\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},u_{k}=1]$. From the properties of expectations, we have the following relation
$\operatorname{\mathsf{E}}\Big[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}\big|\mathcal{I}^{e}_{k},u_{k}\Big]$ $=\operatorname{\mathsf{E}}\Big[\tilde{e}_{k+1}\big|\mathcal{I}^{e}_{k},u_{k}\Big]^{T}\operatorname{\mathsf{E}}\Big[\tilde{e}_{k+1}\big|\mathcal{I}^{e}_{k},u_{k}\Big]$
$\quad+\operatorname{tr}\operatorname{\mathsf{cov}}\Big[\tilde{e}_{k+1}\big|\mathcal{I}^{e}_{k},u_{k}\Big].$
@@
$\tilde{e}_{k+1}$ $=(1-u_{k}\gamma_{k})A_{k}\tilde{e}_{k}+K_{k+1}\nu_{k+1}.$
Accordingly, the conditional mean and the conditional covariance of $\tilde{e}_{k+1}$ given $\mathcal{I}^{e}_{k}$ and $u_{k}$ are obtained as
$\operatorname{\mathsf{E}}\Big[\tilde{e}_{k+1}\big|\mathcal{I}^{e}_{k},u_{k}\Big]$ $=(1-\lambda^{c}u_{k})A_{k}\breve{e}_{k},$
$\operatorname{\mathsf{cov}}\Big[\tilde{e}_{k+1}\big|\mathcal{I}^{e}_{k},u_{k}\Big]$ $=(\lambda^{c}u_{k}-{\lambda^{c}}^{2}u_{k})A_{k}\breve{e}_{k}\breve{e}_{k}^{T}A_{k}^{T}$
$\quad+(1-\lambda^{c}u_{k})A_{k}R_{k}A_{k}^{T}$
$\quad+K_{k+1}N_{k+1}K_{k+1}^{T},$
@@
$\operatorname{\mathsf{E}}\Big[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}\big|\mathcal{I}^{e}_{k},u_{k}\Big]$ $=(1-\lambda^{c}u_{k})\breve{e}_{k}^{T}A_{k}^{T}A_{k}\breve{e}_{k}$
$\quad+(1-\lambda^{c}u_{k})\operatorname{tr}(A_{k}R_{k}A_{k}^{T})$
@@
This implies that $\operatorname{\mathsf{E}}[\tilde{e}_{k+1}^{T}\tilde{e}_{k+1}|\mathcal{I}^{e}_{k},u_{k}]$ is a symmetric function of $\breve{e}_{k}$ and a function of $R_{k}$, and that
@@
$\breve{e}_{k+1}$ $=(1-\bar{\theta}_{k})A_{k}\breve{e}_{k}+K_{k+1}\nu_{k+1},$ (89)
$R_{k+1}$ $=(1-\bar{\theta}_{k})A_{k}R_{k}A_{k}^{T}+\bar{\bar{\theta}}_{k}A_{k}\breve{e}_{k}\breve{e}_{k}^{T}A_{k}^{T},$ (90)
where $\bar{\theta}_{k}=\theta_{k}u_{k}\gamma_{k}+\lambda^{c}(1-\theta_{k})u_{k}$ and $\bar{\bar{\theta}}_{k}=(\lambda^{c}-{\lambda^{c}}^{2})(1-\theta_{k})u_{k}$. By the hypothesis, $V_{k+1}(\breve{e}_{k+1},R_{k+1})$ is a symmetric function of $\breve{e}_{k+1}$ and a function of $R_{k+1}$. Plugging (89) and (90) into $V_{k+1}(\breve{e}_{k+1},R_{k+1})$, we can calculate $\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},u_{k}]$ while all $\nu_{k+1}$, $\theta_{k}$, and $\gamma_{k}$ are averaged out. A straightforward consequence is that we can express $\operatorname{\mathsf{E}}[V_{k+1}(\breve{e}_{k+1},R_{k+1})|\mathcal{I}^{e}_{k},u_{k}]$ as a symmetric function of $\breve{e}_{k}$ and a function of $R_{k}$, since $\nu_{k+1}$ is a Gaussian variable with zero mean. Therefore, we deduce that $\chi^{2}_{k}$ is also a symmetric function of $\breve{e}_{k}$ and a function $R_{k}$. Thus, using (88), we conclude that the claim holds. ∎